i have the following equation where i have to find the value of x, and h & c are the same value
hx-4 = cx/3
so far i did: multiplied the whole equation by 3
3(hx-4)=cx
3hx-12=cx
3hx=cx+12
x=(cx+12)/3h
Could you please help me to get to the answer as i am so stuck
tnx

Answer by Edwin McCravy(20060)

i have the following equation where i have to find 
the value of x, and h & c are the same value

   hx - 4 = cx/3

so far i did: multiplied the whole equation by 3

3(hx - 4) = cx
 3hx - 12 = cx
      3hx = cx + 12

You are correct to here. But you are not ready to
divide yet!  You must first get all terms that 
contain x on the left side of the equation and all
other terms on the right.  So 
subtract cx from both sides:

 3hx - cx = 12

Factor out x:

x(3h - c) = 12

Now you are ready to divide both sides by
(3h - c)

        x = 12/(3h - c)

Edwin